Suppose you expect International Computers Inc (ICI) to pay a dividend of $10 per share in
one year and to trade for $120 per share just after that dividend payment. If ICI is currently
trading at $100 per share and investments with equivalent risk to ICI’s stock have an
expected return of 30%, the capital gain rate is closest to?
Answer #1

CAPITAL GAIN RATE =(P1-P0)/P0

P0 = 100, P1 = 120

CAPITAL GAIN RATE =(P1-P0)/P0 = (120-100)/100 = 20%

Answer : 20% (Thumbs up please)